The error for y is .0078, and the error for x is 0.54. I
am not sure the error for x is correct. x is 1/T and the error of T is 0.54. I am having a hard time finding the error of the slope of the following equation. I am getting a very large error due to the small value of x and I am not sure why. Can someone please help

m =ya-Y Xa - X, = -0.01097-(-1.264) 0.002687 -0.003277k =-2123K
Answer #1

Errortor y = 0.0078 Error for x = 0.54 x = - and T = 0.54 * 1 X= - = . 0.54 M = Ya-Yi - Xarx, = -0.01097-61.1264) 0.00268,15